OpenStudy (anonymous):

identify the solutions of the system of equations if any to -4x - 2y = 12 4y = 8x - 24

OpenStudy (anonymous):

-3 and 0 if that makes since

OpenStudy (anonymous):

4x - 2y = 12 reduces to 2x - y = 6 4y = 8x - 24 -8x + 4y = -24 8x - 4y = 24 reduces to 2x - y = 6 these are the same line, therefore there are infinite solutions